# Sequential Randomization Tests Using E-values: A Betting Approach for Clinical Trials

### v1

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism. Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. The e-RT provides a conservative, assumption-free complement to model-based sequential analyzes.
@@ 1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guaranties may erode.
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guaranty holds at any stopping time, regardless of the stopping rule.
@@ 2.1 Setup
The null hypothesis is that treatment has no effect on any patient. This is the sharp null: every patient would have had the same outcome regardless of assignment. Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
@@ 2.3 Adaptive Betting Strategy
    A patient is randomized (assignment blinded)
@@
    You check the historical records and see in which arm this outcome is more likely
@@
    Wealth updated. If guess is correct, it grows; if wrong, it shrinks
Under the null hypothesis, any historical imbalance between arms is noise. Your bets are sometimes right, sometimes wrong, and wealth fluctuates randomly around 1. Under the alternative, the imbalance is real signal. Your bets are systematically correct, and wealth grows over time.
@@
The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event. The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{1}$. This factor is simply a factor that avoids us from making full bets early on. It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on.
@@ 3.3 Trajectory Examples
Figure 1: e-RT trajectories under the null hypothesis. Panel A: n = 712 (80% power design). Panel B: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.
@@ 4 Discussion
We have presented the e-RT, a nonparametric sequential test for randomized trials based on the betting framework for e-values. The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses.
@@ 4.1 Operating characteristics
Our simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
These results should be interpreted carefully. The e-RT is not designed to replace the primary analysis of a trial; at least not yet. Rather, it provides a continuous monitoring option that requires no alpha spending and no pre-specified interim analysis schedule. If the e-RT crosses its threshold, one could consider stopping the trial. If it does not cross, the trial may proceeds to its planned conclusion and primary analysis.
@@ 4.3 Limitations
There are other limitations. The e-RT tests only whether event rates differ between arms. It does not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used a specific betting strategy (burn-in = 50, ramp = 100); other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is too under development.

### v2

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism. Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. The e-RT provides a conservative, assumption-free complement to model-based sequential analyses.
@@ 1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guarantees may erode.
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
@@ 2.1 Setup
The null hypothesis is that treatment has no effect. Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
@@ 2.3 Adaptive Betting Strategy
    A patient is randomized (assignment blinded).
@@
    Historical records are checked. The most likely arm for this outcome is defined.
@@
    Wealth updated. If guess is correct, it grows; if wrong, it shrinks.
Under the null hypothesis, any historical imbalance between arms is noise. Bets will be sometimes right, sometimes wrong, and wealth fluctuates randomly around 1. Under the alternative, the imbalance is real signal. Bets will be systematically correct, and wealth grows over time.
@@
The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event. The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{i}$. This factor is simply a factor that avoids us from making full bets early on. It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on. Finally, the $0.5$ multiplier before $c_{i}$ and the wager ($\hat{\delta}_{i-1}$) simply keeps $\lambda$ between zero and one.
@@ 2.4 Worked Example
In summary: the outcome determines $\lambda$ (which direction to bet); the treatment arm determines which multiplier formula to apply (whether the bet paid off). The machinery is agnostic to whether the effect size is following the expected direction; it simply keeps betting according to what is more likely.
@@ 3.3 Trajectory Examples
Figure 1: e-RT trajectories under the null hypothesis. Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.
@@ 4 Discussion
The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values. The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses.
@@ 4.1 Operating characteristics
Simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
These results should be interpreted carefully. The e-RT is not designed to replace the primary analysis of a trial; at least not yet. Rather, it provides a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule. If the e-RT crosses its threshold, one could consider stopping the trial. If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
@@ 4.3 Limitations
There are other limitations. The e-RT tests only whether event rates differ between arms. It does not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used a specific betting strategy (burn-in = 50, ramp = 100); other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is also under development.
@@ 4.5 Disclaimer
4.6 Version Control
  1. 1.
    First Version (Dec 04, 2025)
  2. 2.
    Second Version (this version): Minor text adjustments; removed claim on sharp null.